OpenStudy (anonymous):

an initial investment of $1,500 over 25 years at 5% APR. After 10 years about how much interest has the investment earned?

OpenStudy (allank):

Assuming that continuous compounding is in effect, the investment value can be tracked exponentially by the expression: v(t)=1500e^(r*t) where r=0.05 (rate per year) and t is the number of years Plug in any value of t and you'll know the investment value after those number of years have elapsed

ganeshie8 (ganeshie8):

problem needs to specify whether it is simple interest or compoinded yearly @SerinaRenee

OpenStudy (anonymous):

compounded

ganeshie8 (ganeshie8):

compounded yearly, right ?

ganeshie8 (ganeshie8):

if they dont specify the period of compounding, we can assume it is yearly as it is the case for most banks..

ganeshie8 (ganeshie8):

check top of your question and see if they tell anything about period of compounding

OpenStudy (anonymous):

yeah its yearly

ganeshie8 (ganeshie8):

great ! we can use compound interest formula then... can you check your notes and see if you have it handy

OpenStudy (anonymous):

A=P(1+r)^t

ganeshie8 (ganeshie8):

ty :) lets plugin the values

ganeshie8 (ganeshie8):

P = initial investment = 1500 r = interest rate = 5 % = 5/100 t = time = 10 A = final investment

ganeshie8 (ganeshie8):

plug them and see what you get for final investment A

OpenStudy (kenljw):

Do you guy's actually remember formula first year 1,500(1+.05) second year 1,500(1+.05)(1+.05)=1,500(1+.05)^2 24 years 1,1500(1+.05)^25=5,079.53 I can never remember formulas that why I leaned to derive them